OpenStudy (anonymous):

Write an expression(fully simplified)that represents the area of a square that has a side length of 2x+12 units

OpenStudy (anonymous):

A = w * h A = (2x + 12)(2x+12) A = (2x+12)^2

OpenStudy (anonymous):

tks again fruitless your a life saver